Let A denote the event 'student is female' and let B denote the event 'student is
French'. In a class of 100 students suppose 60 are French, and suppose
that 10 of the French students are females. Find the probability that if I pick a
French student, it will be a girl that is, find P (A|B).
1
Expert's answer
2020-09-28T19:17:36-0400

Class has 100 students.

Probability to pick up French is "\displaystyle P(B) =\frac{60}{100}= 0.6" .

Probability to pick up female French is "\displaystyle P(A\cap B) = \frac{10}{100} = 0.1".

"\displaystyle P(A|B) = \frac{P(A\cap B)}{P(B)} = \frac{0.1}{0.6}= \frac{1}{6}"

Answer: 1/6.